OpenStudy (anonymous):

find the solution set: t(3t + 1) = 0

OpenStudy (unklerhaukus):

if a product is equal to zero one of the terms must be equal to zero t=0 or 3t+1=0

OpenStudy (anonymous):

t(3t+1)=0 either t=0 or 3t+1=0 so t=0,(-1/3)

OpenStudy (anonymous):

t=0

OpenStudy (anonymous):

thanks I knew it was 0 but I thought there was something I was missing

OpenStudy (unklerhaukus):

can you see how to get the second solution/

OpenStudy (anonymous):

yes thanks, that was what I was looking for